The probability that a couple will have n baby boys in a row is given by the formula (1/2)to the n. Find the probability that a couple will have four baby boys in a row.
Found 2 solutions by solver91311, richard1234:
Answer by solver91311(24713)





Just do the arithmetic.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by richard1234(7193)

P(4 in a row) = (1/2)^4 = 1/16